A 30 ft by 40 ft rectangle has two squares of side x removed from it. The remaining area is 1000 square feet. Find x
Answer by mananth(16946)

30*40 = 1200 sq.ft
...
side x of square
area = x^2
2 squares = 2x^2
...
1200-2x^2= 1000
2x^2=200
/2
x^2= 100
take squareroot
x= 10 ft
...
m.ananth@hotmail.ca